OpenStudy (anonymous):

Minnie deposited $5,375 into a savings account 18 years ago. The account has an interest rate of 3.2% and the balance is currently $9,561.39. How often does the interest compound?

OpenStudy (tkhunny):

What have you tried?

OpenStudy (anonymous):

I tried it using the compound formula A=P(1+r/n)^nt and got the answer wrong. Am I using the wrong formula? @tkhunny

OpenStudy (anonymous):

Ift it compounds annually, I got $9476 from your formula, close. You could try some other common compounding practices, such as twice a year, which gave $9518. Getting loser.

OpenStudy (anonymous):

Monthly?

OpenStudy (anonymous):

I calculated it quarterly and got an almost similar answer. I got $9,539.55 . Can you do it and see if you got the same answer as me? @douglaswinslowcooper

OpenStudy (anonymous):

I got 9539.74, almost identical. Go to 6 times a year or better 12 times a year.

OpenStudy (anonymous):

12 per year gave me 9554. Radical move: 365 times a year, daily, Bingo: $9561

OpenStudy (anonymous):

There is an exponential function for truly continuous compounding, exp(something). That may be what was sought. We got close.

OpenStudy (anonymous):

A/P = exp[(0.032)(t)] That gives $9561.63, which may be what they wanted you to do.

OpenStudy (anonymous):

Good night.

OpenStudy (tkhunny):

A trip down logical paths, using existing compounding periods, may solve the problem. First, let's see if it's even possible. Annual -- n = 1 -- $9,475.74 Lowest Possible Continuous -- n = \(\infty\) -- $9,561.63 Highest Possible Good, the desired answer is between those. It's pretty close to the Continuous, so I suspect it will require rather frequent compounding. Semi-Annual -- n = 2 -- $9,518.14 Quarterly -- n = 4 -- $9,539.75 Monthly -- n = 12 -- $9,554.31 Weekly -- n = 52 -- $9,559.94 Daily -- n = 365 -- $9,561.39 There it is!